In a survey of 500 infants chosen at random, it is found that 240 are girls. Are boy and girl births equally likely according to this survey (use α = 0.05)


1
Expert's answer
2021-07-14T18:15:27-0400
"\hat{p}=\dfrac{240}{500}=0.48"

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p=0.5"

"H_1:p\not=0.5"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test is "z_c=1.96."

The rejection region for this two-tailed test is "R=\{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.48-0.5}{\sqrt{\dfrac{0.5(1-0.5)}{500}}}=-0.8944"

Since it is observed that "|z|=0.8944<1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value is "p=2P(Z<-0.8944)=0.3711," and since "p=0.3711>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is different than "0.5," at the "\alpha=0.05" significance level.

Therefore, there is enough evidence to claim that boy and girl births are equally likely, at the "\alpha=0.05" significance level.